10.A cars suspension can be modeled by 5000x + 30000x + 45000x = 0 . Find the solution assuming that x(0)=9 x(0)= 2 . Gra
Answer #1

Given sooo -5000-12 + 30000 x +4 5000x = 0 +6*-+-92) = 0 11 x + 6x + 9x = 0 Corresponding Auxillary Equation is D+60 + 9 =* only one times Graph Goss the t-axis at E = 9 29. * Because at t= -9 29 Yog 29